Sample of glucose C6H12O6 contains 1.250x10^21 atoms of carbon.
How many atoms of hydrogen does it contain?
How many molecules of glucose does it contain? How many moles of glucose does it have?
What's the mass of this sample in grams?
1
Expert's answer
2015-06-08T07:34:38-0400
Each molecule of glucose contains two atoms of Hydrogen per one atom of Carbon. So, the sample contains 2 x 1.25 x 1021 = 2.5x1021 atoms of Hydrogen.
6 atoms of Carbon - 1 molecule of glucose
1.250 x 1021 atoms of Carbon - X molecules,
So, the sample contains X = 1.250x1021/6 = 0.208x1021 molecules of glucose
1 mole of glucose - 6.022x1023 molecules
n mole of glucose - 0.208x1021 molecules, 
So, the sample contains n = 0.208x1021/6.022x1023 = 3.45x10^4 moles of glucose
m=Mxn, where M - molar mass of glucose (180 g/mol), n - moles
So, the massm of the sample is m = 180 x 3.45x10^24 = 0.62 g